### v1

# Minimal thermodynamic cost of computing with circuits

Abstract
All digital devices have components that implement Boolean functions, mapping that component’s input to its output. However, any fixed Boolean function can be implemented by an infinite number of circuits, all of which vary in their resource costs. This has given rise to the field of circuit complexity theory, which studies the minimal resource cost to implement a given Boolean function with any circuit. Traditionally, circuit complexity theory has focused on the resource costs of a circuit’s size (its number of gates) and depth (the longest path length from the circuit’s input to its output). In this paper, we extend circuit complexity theory by investigating the minimal thermodynamic cost of a circuit’s operation. We do this by using the mismatch cost of a given circuit that is run multiple times in a row to calculate a lower bound on the entropy production (EP) incurred in each such run of the circuit. Specifically, we derive conditions for mismatch cost to be proportional to the size of a circuit, and conditions for them to diverge. We also use our results to compare the thermodynamic costs of different circuit families implementing the same family of Boolean functions. In addition we analyze how heterogeneity in the underlying physical processes implementing the gates in a circuit influences the minimal thermodynamic cost of the overall circuit. These and other results of ours lay the foundation for extending circuit complexity theory to include mismatch cost as a resource cost.
I Introduction
Computational complexity theory explores the relative difficulty of computing functions within formal models of computation. In his foundational work turing1936computable , Alan Turing introduced the concept of Turing machines and demonstrated that functions can be classified as either computable or non-computable. However, even within the class of computable functions, some are inherently harder to compute than others rabin1960degree ; cook2007overview . The fundamental question arises: what does it mean for one function to be more difficult to compute than another?
A given function can often be computed using multiple algorithms, each incurring different resource costs depending on the model of computation. Traditionally, computational complexity theory has focused on quantifying resource usage such as the number of computational steps or memory requirements. Early discussions on complexity measures also considered physical energy expenditure and thermodynamic work as potential complexity metrics cobham1965intrinsic .
Landauer and Bennett were among the first to explore the thermodynamic cost of computation using equilibrium statistical mechanics. They famously argued that erasing a single bit of information necessarily generates at least $k_{B}\ln 2$ of heat, where $k_{B}$ is Boltzmann’s constant and $T$ is the temperature of the surrounding heat bath landauer1961irreversibility ; bennett1982thermodynamics . However, this early work modeled computation—which is inherently a non-equilibrium process—within the framework of equilibrium thermodynamics. As a result, the analysis was semi-formal and limited in scope, leaving many fundamental questions unanswered. At the time, the theoretical tools needed to rigorously analyze far-from-equilibrium systems were not yet available, and the thermodynamics of computation remained an open and underdeveloped area.
Fortunately, recent advances in non-equilibrium thermodynamics, particularly within the framework of stochastic thermodynamics, have made it possible to rigorously define and analyze thermodynamic quantities such as work, entropy, and heat dissipation in systems driven far from equilibrium van2015ensemble ; seifert2012stochastic . These tools are now being applied to fundamental questions about the thermodynamic costs of computation and communication processes van2015ensemble ; seifert2012stochastic ; wolpert2019stochastic ; wolpert2024thermodynamics ; wolpert2024stochastic . In particular, stochastic thermodynamics allows to analyze the entropy production (EP) of running any (physical system that implements) a computation, i.e., the unavoidably wasted energy incurred by running that computation.
Boolean circuits provide an alternative model of computation, characterizing the complexity of a function in terms of the size and depth of the circuit implementing it. The size of a circuit is the total number of gates, while the depth corresponds to the longest path from an input to an output, representing the time required for computation arora2009computational ; vollmer1999introduction . Just as there exist multiple algorithms for a given problem, there are infinitely many Boolean circuits that can implement a given Boolean function. The task of identifying the most efficient circuit among those infinitely many circuits is both theoretically difficult and practically important. Circuit complexity theory provides a framework for comparing and optimizing circuits based on size and depth, but the associated thermodynamic cost of running a circuit as a key measure in optimization has not been given considerable attention.
As processors shrink toward their theoretical efficiency limits—and as computational demands surge, especially due to large-scale data processing—understanding the energy cost of computation is becoming increasingly important. Incorporating thermodynamic considerations into circuit complexity theory offers a principled framework to quantify minimal heat dissipation. This perspective can refine existing complexity measures and inform the design of circuits that are both logically and thermodynamically efficient. Moreover, since circuits determine whether a function is efficiently parallelizable or inherently sequential, thermodynamic analysis can illuminate the energy efficiency of parallel computation.
A central concept in this analysis is the mismatch cost (MMC) kolchinsky2021dependence ; kolchinsky2017dependence , which provides a lower bound on entropy production in physical computational processes.
Figure 1: An example of a 2-input and 1-output circuit with a topological ordering.
Contributions and Roadmap
This paper presents several key contributions aimed at understanding the thermodynamic cost of computation in Boolean circuits. First, we derive a lower bound on the EP arising in running Boolean circuits. This bound emerges primarily from the modular structure of circuits, where the dynamics of gates depend on the states of only a subset of gates, creating a network of dependencies that govern the flow of logic and dynamics. Based on this bound, we introduce a new circuit complexity measure—mismatch cost complexity. We also establish formal relationships between MMC complexity and traditional size complexity. Additionally, we compute the MMC complexity for various circuit families that implement the same Boolean function, enabling a comparison of their thermodynamic efficiency.
The structure of the paper is as follows: In Sec. II.1, we provide a minimal background on circuit complexity theory, reviewing the concepts of circuit families, size and depth complexity, and their associated complexity classes. In Sec. II.2, we briefly review stochastic thermodynamics and the mismatch cost lower bound on EP. In Sec. III, we focus on computing the mismatch cost for circuits. The circuit diagram alone does not specify the dynamical changes of the circuit’s state, i.e., of the joint distribution over the circuits gates as it runs. In Sec. III.1 we explain what other information is needed, and how to use it to calculate that change in the state of a circuit evolves during its execution. In Sec. III.2 and Sec. III.3 we combine these results, to derive the expression for the total MMC in computation with circuits.
Sec. IV builds on this earlier analysis to introduce mismatch cost complexity as a new circuit complexity measure, analogous to size and depth complexity, representing the minimal energetic cost of implementing a Boolean function with any circuit. We derive results and theorems that link MMC complexity to size complexity, showing that under certain conditions, the upper bound on MMC grows linearly with the size complexity of a circuit family. Complementing this result, we identify conditions where MMC scales differently from size complexity. In Sec.V, we discuss the energetic aspects of circuit optimization, comparing circuit families that implement a given Boolean function. Specifically, we study two families that compute the ADD function—one with depth complexity $\log(n)$ and another with $n$—to compare their mismatch costs. We also investigate the impact of heterogeneity in the physical properties of gates on the associated energetic costs. Finally, Sec. VI concludes the paper by discussing possible directions for future research.
II Background
II.1 Circuit complexity theory
A circuit is built from basic logic gates that are interconnected to process inputs and produce outputs. A basis is a set $\mathcal{B}$ of Boolean functions that define the allowable gates in a circuit. For example, a basis containing AND and NOT gates is considered a universal basis because any Boolean function can be implemented using only these gates. Formally, a Boolean circuit is represented as a directed acyclic graph (DAG), denoted by $(V,E,\mathcal{B},\mathcal{X})$. Here, $V$ is the set of nodes, where each node represents a logic gate selected from a finite basis $\mathcal{B}$. The edges in $E$ define the directed connections between gates, determining how information flows through the circuit. The set $\mathcal{X}$ represents the state space of the circuit (see below). The direction of edges reflects dependencies between gates, enabling a specific ordering of nodes in $V$. This ordering, called a topological ordering, ensures that for every directed edge from node $\mu$ to node $\nu$, node $\mu$ appears before node $\nu$ in the sequence ($\mu<\nu$).
A node $\mu$ has incoming edges from nodes known as its parent nodes, denoted as $\mathrm{pa}(\mu)$, and outgoing edges to nodes known as its children nodes, denoted as $\mathrm{ch}(\mu)$. Nodes that have no incoming edges, called input nodes or root nodes, serve as the circuit’s inputs. We denote the set of input nodes as $V_{in}\subset V$ and the set of all non-input nodes as $V_{nin}=V\setminus V_{in}$.
Each gate $\mu$ in the circuit has an associated state space, denoted by $X_{\mu}$, and its state is represented by $\boldsymbol{x}_{\mu}\in X_{\mu}$. For binary gates, the state space is $X_{\mu}=\{0,1\}$. The collective states of all gates together define the joint state of the circuit. The joint state space of the circuit is given by
$\mathcal{X}=\bigotimes_{\mu}X_{\mu}$
and the joint state of the circuit is denoted by $\boldsymbol{x}$.
We use $\boldsymbol{x}_{\mathrm{pa}(\mu)}$ and $\boldsymbol{x}_{\mathrm{ch}(\mu)}$ to denote the joint state of the parent and children gates of a gate $\mu$, respectively. Similarly, we define $\boldsymbol{x}_{in}$ and $\boldsymbol{x}_{nin}$ to represent the joint states of all input nodes and all non-input nodes, respectively.
The size of the circuit is defined as the total number of gates, or $|V|$. The depth of the circuit is defined as the length of the longest path from an input node to an output node. Circuit size and depth are two important complexity measures of a circuit, they also provide a measure of the hardness of computation of a function $f$ as the size and depth of the minimal circuit that computes it.
An $n$-input Boolean function is a mapping
$f^{n}:\{0,1\}^{n}\to\{0,1\}^{m}$
which takes $n$-bit binary inputs and produces $m$-bit binary outputs. A family of Boolean functions is defined as a sequence $f=\{f^{n}\}_{n\in\mathbb{N}}$, where each $f^{n}$ corresponds to a specific input size $n$. This formulation allows us to define an associated function
$f:\{0,1\}^{*}\to\{0,1\}^{*}$
that operates on inputs of arbitrary length. For example, consider the addition function
$\text{ADD}^{n}:\{0,1\}^{2n}\to\{0,1\}^{n+1}$
which takes two $n$-bit binary representations of natural numbers and outputs their sum in binary. The function family $\{\text{ADD}^{n}\}_{n\in\mathbb{N}}$ extends this addition operation to all possible input sizes.
Circuits differ fundamentally from other models of computation. In models like Turing machines, a single machine accepts input of of any arbitrary length. In contrast, circuits require a distinct specification for each input length. This means that to compute functions belonging to the same family but with different input sizes, one must design distinct circuits for each case. Because the model does not provide a uniform mechanism for handling all input lengths, it is classified as a non-uniform model of computation arora2009computational ; vollmer1999introduction . As a result, computation in this model is described using a circuit family, denoted as $\{C_{n}\}_{n\in\mathbb{N}}$, where each circuit $C_{n}$ corresponds to inputs of length $n$. A circuit family $\mathcal{C}~{}$ is a sequence $\{C_{n}\}_{n\in\mathbb{N}}$ of Boolean circuits, where $C_{n}$ is the circuit with $n$ inputs.
A circuit family $\mathcal{C}$ is said to compute a function $f$ if, for every input string $w$ of length $n$, the circuit correctly evaluates the function:
$f^{n}(w)=C_{n}(w)$
for all $n\in\mathbb{N}$. The size complexity of a circuit family is the number of gates in $n$ as a function of $n$, while the depth complexity is the length of the longest directed path from an input node to an output node, also expressed as a function of $n$.
Definition 1 (Size and depth complexity)
A circuit family of size complexity $s(n)$, for a function $s:\mathbb{N}\to\mathbb{N}$, is any circuit family where the size $|C_{n}|\leq s(n)$ for every $n$.
Analogously, a circuit family of depth complexity $d(n)$, for a function $d:\mathbb{N}\to\mathbb{N}$, is any circuit family where the depth of $C_{n}\leq d(n)$ for every $n$.
Likewise, one can define the size and depth complexity of a function $f$ as the size and depth complexity of the associated circuit family that computes it.
Definition 2 (Circuit complexity class)
Let $s:\mathbb{N}\to\mathbb{N}$. $\mathbf{SIZE}(s)$ is the class of functions $f$ for which there exist a circuit family $\mathcal{C}~{}$ of size complexity $s$.
Analogously, let $d:\mathbb{N}\to\mathbb{N}$. $0pt{d}$ is the class of functions $f$ for which there exist a circuit family $\mathcal{C}~{}$ of depth complexity $d$.
The growth of a circuit’s size or depth with respect to the input length of a function provides a crucial measure of the function’s computational complexity. In his seminal work, Shannon proved that most Boolean functions require circuits of exponential size, specifically showing that there exist functions that cannot be computed by circuits smaller than $2^{n}/10n$. In other words, the vast majority of functions—often referred to as "hard" functions—such that $f\notin\mathbf{SIZE}(2^{n}/10n)$ and thus require exponentially large circuits to compute shannon1949synthesis . However, despite this result, explicitly identifying such a function has remained challenging, as most functions of practical interest can still be computed by circuits of reasonable, albeit large, size.
An interesting class of functions emerges when restricted to "small" circuits, i.e., circuit families whose size does not grow faster than polynomial with input length. $\mathbf{P_{/poly}}$ is the class of functions for which there exist a $\mathrm{poly}(n)$-size circuit family,
$\mathbf{P_{/poly}}:=\cup_{c\geq 0}\mathbf{SIZE}(n^{c}).$
The class $\mathbf{NC}^{d}$ consists of functions that can be computed by a circuit family of polynomial size $\mathrm{poly}(n)$ and depth at most $\log^{d}n$, for some $d\in\mathbb{N}$. Since the depth, which represents the maximum number of computation steps, grows at most as a logarithmic function of the input size raised to the power of $d$, the functions in $\mathbf{NC}^{d}$ can be computed very quickly. These classes form the $\mathbf{NC}$ hierarchy, which is defined to capture the notion of problems that have very fast parallel algorithms using a feasible amount of hardware.
$\mathbf{NC}=\bigcup_{d}\mathbf{NC}^{d}.$
A function has an efficiently parallel algorithm if and only if it is in $\mathbf{NC}$ arora2009computational ; vollmer1999introduction .
The previous example of ADD is in $\mathbf{NC}$. However, just like any Boolean function, ADD has many circuit families that can implement it. For example, ADD can be implemented using two well-known circuit families: the ripple-carry adder (RCA) and the carry look-ahead adder (CLA), each offering distinct trade-offs in size and depth complexity vollmer1999introduction .
The RCA follows the simple "carry-over" method taught in high school arithmetic, where each bit addition depends on the carry from the previous bit. As a result, for an $n$-bit input, the depth of the circuit grows linearly with $n$. Additionally, increasing the input size by one bit requires adding one additional full-adder circuit to the chain. This means that the overall circuit size also grows linearly with $n$. Its simplicity makes it easy to scale. However, since each stage of bit addition must wait for the carry from the previous stage before proceeding, RCA becomes significantly slower as the input size increases.
In contrast, CLA uses a more sophisticated approach by predicting carry bits in advance rather than waiting for them to propagate sequentially. This significantly reduces the depth of the circuit, which grows proportionally to the logarithm of the input size. As a result, computation time decreases drastically for large inputs. However, the parallel computation and complex wiring cause the total number of gates in the circuit to grow at a rate proportional to $\log n$. Consequently, as the input size increases, the CLA circuit becomes increasingly intricate.
This exemplifies the trade-offs involved in designing a circuit family to implement a Boolean function. In addition to size and depth, another crucial complexity measure is the energy dissipated in a circuit, which has not yet been considered. In the next section, we briefly introduce stochastic thermodynamics and its framework for computing irreversible heat dissipation in a circuit. This provides an additional complexity measure, analogous to size and depth, that quantifies the energetic cost of a circuit family.
II.2 Stochastic Thermodynamics and Mismatch Cost
Real-world computational systems operate far from equilibrium, with multiple interdependent components evolving rapidly. Traditional equilibrium statistical mechanics, which applies only to large macroscopic systems that either evolve quasi-statically or remain static, is inadequate for analyzing the energetic costs of computation wolpert2024stochastic . Mesoscopic systems far from equilibrium that exhibit stochastic fluctuations can be analyzed within the framework of stochastic thermodynamics. The key assumption in this framework is that any degrees of freedom not explicitly described by the system’s dynamics—such as internal states or the reservoirs—remain in equilibrium. This assumption allows one to relate thermodynamic quantities to system dynamics through the principle of local detailed balance seifert2012stochastic ; seifert2018stochastic ; esposito2010entropy ; esposito2012stochastic .
One of the central quantities in stochastic thermodynamics is entropy production (EP), which quantifies the irreversible heat dissipation in a process. If the probability distribution over the states of a physical system during its evolution changes from $p_{0}$ to $p_{\tau}$, the total EP associated with an initial distribution $p_{0}$ can be expressed as van2015ensemble ; kolchinsky2017dependence ,
$\sigma(p_{0})=Q(p_{0})-[S(p_{\tau})-S(p_{0})],$ (1)
where $Q(p_{0})$ is called the total entropy flow and it corresponds to the change in the thermodynamic entropy of the environments during the process. If the system interacts with $N$ heat reservoirs, the total entropy flow is given by the sum of the energy exchanged with each reservoir, weighted by the inverse of its temperature:
$Q=\sum_{i=1}^{N}\frac{\mathcal{Q}_{i}}{T_{i}},$ (2)
where $\mathcal{Q}_{i}$ represents the net energy flow into $i$-th reservoir at temperature $T_{i}$, and $k_{B}$ is Boltzmann’s constant. In the context of computation, let $X$ represents the logical states of a computer and $G$ describes a computational map such that after the end of computation, an initial distribution $p_{0}$ evolves to $p_{\tau}(x^{\prime})=\sum_{x}G(x^{\prime}|x)p_{0}(x)$, or as a shorthand $p_{\tau}=Gp_{0}$.
Advances in stochastic thermodynamics have led to the discovery of various contributions to total EP based on key properties of a system’s dynamics. Notably, thermodynamic uncertainty relations (TURs) establish a contribution to EP that arises from he precision of currents involved in the system’s evolution barato2015thermodynamic ; gingrich2016dissipation ; horowitz2020thermodynamic . Another significant class of results, known as speed limit theorems (SLTs), provide a lower bound on EP in terms of the average number of state transitions during the system’s evolution shiraishi2018speed ; vo2020unified . In the context of computation, these results depend on the nitty-gritty details of how the map $G$ is implemented—for instance, SLTs require knowledge of the average number of transitions in the underlying dynamics, while TURs depend on the precision of currents involved in the process.
Another key contribution to EP is the mismatch cost. Consider the initial distribution that minimizes EP in Eq. 1, denoted as $q_{0}$, which we refer to as the prior distribution of the process. When the process starts from a different initial distribution $p_{0}$, the resulting EP, $\sigma(p_{0})$, can be decomposed as kolchinsky2017dependence ; kolchinsky2021dependence ,
$\sigma(p_{0})=[D(p_{0}||q_{0})-D(Gp_{0}||Gq_{0})]+\sigma_{\text{res}},$ (3)
where $D(p_{0}||q_{0})$ is the Kullback-Leibler (KL) divergence between $p_{0}$ and $q_{0}$, and $\sigma_{\text{res}}$, known as the residual EP, represents the minimum EP of the process, defined as $\sigma_{\text{res}}:=\sigma(q_{0})$. The drop in KL-divergence $D(p_{0}||q_{0})-D(Gp_{0}||Gq_{0})$ is called the mismatch cost. The MMC is always non-negative, a consequence of the data processing inequality for KL-divergence. Unlike the residual EP $\sigma_{\text{res}}$, which depends on the fine details of the physical implementation, MMC depends only on the computational map $G$, the initial distribution $p_{0}$, and the prior distribution $q_{0}$. Additionally, $\sigma_{\text{res}}$ is non-negative due to the second law. In certain dynamical situations, the MMC provides a strictly positive contribution to EP. Consider a process where the computational map $G$ is repeatedly applied to the state space. Since the underlying physical process implementing $G$ remains unchanged across iterations, the prior distribution $q_{0}$ also remains the same. However, the actual state distribution evolves with each iteration: after the $i$-th iteration, it is given by $p_{i}=Gp_{i-1}$, or more generally, $p_{i}=G^{i}p_{0}$. The total MMC then accumulates over iterations as:
$\mathcal{MC}(p_{0})=\sum_{i=0}^{\tau}D\left(G^{i}p_{0}||q_{0}\right)-D\left(G^{i+1}p_{0}||Gq_{0}\right)$ (4)
Even if the process starts at the prior $q_{0}$, after the first iteration, the distribution becomes $p_{1}=Gq_{0}$, which differs from $q_{0}$. This deviation from the prior distribution results in a strictly positive MMC in the next iteration, and the same holds for subsequent iterations wolpert2024thermodynamics .
Importantly, this strictly positive lower bound on EP holds for any initial distribution $p_{0}$, regardless of the prior $q_{0}$. As the system evolves through a sequence of distributions $\{p_{0},p_{1},\dots,p_{\tau}\}$, where $p_{t+1}=Gp_{t}$, there exists a distribution $\hat{q}_{0}$ that minimizes the expression on the right-hand side of Eq. 4. The associated MMC, denoted as $\mathcal{MC}_{\hat{q}_{0}}(p_{0})$, provides a strictly positive lower bound on the actual MMC, regardless of the prior distribution $q_{0}$.
Thus, this strictly positive periodic MMC contribution to EP is entirely independent of the underlying physical process and is fundamentally tied to the computational map $G$ and its repeated application, which highlights the unavoidable EP contribution intrinsic to the computation. In the next section, we describe how the probability distribution over a circuit’s states evolves during execution, and how the mismatch between the evolving distribution and the optimal prior distribution contributes to the EP.
Figure 2: State space dynamics of a circuit computation. (A) The circuit begins in a state inherited from the previous run, with all gate values logically dependent according to the circuit structure. (B) The inputs are updated for the new run, resetting their values independently of the rest of the circuit. (C–G) Gates are sequentially updated based on their input dependencies, progressively building logical correlations as computation unfolds.
III Mismatch cost of computing with circuits
III.1 Dynamics of the circuit
A circuit diagram alone does not specify the dynamics of the joint state of the physical circuit, the order in which the gates are run, and how the state of each gate updates, etc. In this section, we introduce one relatively simple way to specify these details from the circuit diagram.
In this specification, the gates are run serially, one after the other. The topological ordering of a circuit determines the sequence in which its gates are executed. Accordingly, we use the topological index $\mu$ as a step index, meaning that gate $\mu$ updates its value at step $\mu$ (see Fig. 2). Once the input nodes are assigned new values, each gate updates sequentially in the prescribed topological order, with gate $\mu$ updating based on the values of its parent gates. After one complete execution, the output gates provide the final result corresponding to the given input, and moreover, the circuit is fed with new input values for the next run while all gate values remain unchanged. The same sequence of updates is repeated, with gates updating in topological order based on the new inputs and the latest values of their parents.
Due to this repeated use of the circuit, the probability distribution over the state of the circuit goes through a cycle, as depicted in Fig. 2 and Fig. 4 and explained below: after a complete execution of the circuit, the states of the gates are correlated with each other, as depicted in Fig. 2A. Importantly, the non-input gates are correlated with the input nodes. However, when new input values are assigned for the next execution, the input nodes become uncorrelated with the values of the non-input gates from the previous execution (Fig. 2B). Next, as the sequential process begins, and each gate updates its value based on the new values of its parents, there is a making and breaking of correlation. Before gate $\mu$ updates, it remains correlated with its children gates $\mathrm{ch}(\mu)$ from the previous execution. However, since its parent gates $\mathrm{pa}(\mu)$ have already been updated in the current execution, $\mu$ is not correlated with them at this stage. When gate $\mu$ updates, two simultaneous changes occur: first, $\mu$ becomes correlated with its parent gates $\mathrm{pa}(\mu)$ in the current execution, and second, its new value becomes independent of the values of its children gates $\mathrm{ch}(\mu)$ from the previous execution.
We assume input values are jointly sampled from an input distribution $p_{in}(\boldsymbol{x}_{in})$. The dependency of a gate on its parents is expressed by the conditional distribution $\pi_{\mu}(\boldsymbol{x}_{\mu}|\boldsymbol{x}_{\mathrm{pa}(\mu)})$ for each non-input gate $\mu\in V_{nin}$. The conditional distributions $\pi_{\mu}$ together specify the conditional probability of the state of the non-input gates given the states of input gates:
$\mathrm{R}(\boldsymbol{x}_{nin}|\boldsymbol{x}_{in})=\prod_{\mu\in V/V_{in}}\pi_{\mu}(\boldsymbol{x}_{\mu}|\boldsymbol{x}_{\mathrm{pa}(\mu)})$ (5)
Note that, regardless of the gate being deterministic or noisy, since the inputs are randomly sampled from an input distribution $p_{in}$, the state of each gate, and hence the joint state of the circuit is a random variable. After the end of a complete run and before the beginning of the next run, the joint state of the entire circuit has a following probability distribution (see Fig. 3A)
$p^{0}(\boldsymbol{x})=R(\boldsymbol{x}_{nin}|\boldsymbol{x}_{in})p_{in}(\boldsymbol{x}_{in})$ (6)
Table of notation
Symbol Definition
$C_{n}$ Circuit with input size $n$
$\mu$ Index of a gate in circuit, $\mu\in\{1,2,...,|C_{n}|\}$
$\mathrm{pa}(\mu)$ Set of parent gates of gate $\mu$
$\mathrm{ch}(\mu)$ Set of children gates of gate $\mu$
$V$ Set of input nodes and non-input gates in circuit
$V_{l}$ Set of gates in layer $l$ of circuit
$\mathrm{X}_{\mu}$ Random variable representing the states of gate $\mu$
$\mathrm{X}_{l}$ Random variable representing the joint state of gates in layer $l$.
$\mathrm{X}_{\chi}$ Random variable representing the joint state of subset of gates $\chi$.
$\boldsymbol{x}$ Joint state of the entire circuit
$\boldsymbol{x}_{\mu}$ State of gate $\mu$
$\boldsymbol{x}_{\mathrm{pa}(\mu)}$ Joint state of parent gates of $\mu$
$\boldsymbol{x}_{:\mu}$ Joint state of gates preceding gate $\mu$, including input nodes
$\boldsymbol{x}_{l}$ Joint state of gates in layer $l$
$\boldsymbol{x}_{:l}$ Joint state of gates in layers preceding layer $l$
$p^{0}(\boldsymbol{x})$ Maximally correlated distribution over the circuit states at the start of a new run
$p^{\mu}(\boldsymbol{x})$ Distribution over states of circuit before gate $\mu$ runs in a gate-by-gate execution
$p^{l}(\boldsymbol{x})$ Distribution over states of circuit before layer $l$ runs in a layer-by-layer execution
$p_{\chi}(\boldsymbol{x}_{\chi})$ Marginal distribution of subset $\chi$ of gates obtained from $p^{0}(\boldsymbol{x})$
$p_{:\mu}(\boldsymbol{x}_{:\mu})$ Marginal distribution of subset of gates preceding gate $\mu$
$p_{\mu:}(\boldsymbol{x}_{:\mu})$ Marginal distribution of subset of gates following gate $\mu$
$p_{l}(\boldsymbol{x}_{l})$ Marginal distribution of subset of gates in layer $l$ obtained from $p^{0}(\boldsymbol{x})$
$q^{\mu}(\boldsymbol{x})$ Prior distributions of process of updating gate $\mu$ in circuit
$q^{0}(\boldsymbol{x})$ Prior distribution of the process of overwriting the inputs with new values
$\tilde{q}^{\mu}(\boldsymbol{x})$ Distribution obtained by evolving $q^{\mu}(\boldsymbol{x})$ as gate $\mu$ updates
$q_{\mu,\mathrm{pa}(\mu)}(\boldsymbol{x}_{\mu},\boldsymbol{x}_{\mathrm{pa}(\mu)})$ Prior distribution associated with gate $\mu$
$\tilde{q}_{\mu,\mathrm{pa}(\mu)}(\boldsymbol{x}_{\mu},\boldsymbol{x}_{\mathrm{pa}(\mu)})$ Distribution obtained by evolving $q_{\mu,\mathrm{pa}(\mu)}(\boldsymbol{x}_{\mu},\boldsymbol{x}_{\mathrm{pa}(\mu)})$ under the update of gate $\mu$
$q_{\mathrm{pa}(\mu)}(\boldsymbol{x}_{\mathrm{pa}(\mu)})$ Marginal distribution obtained from $q_{\mu,\mathrm{pa}(\mu)}(\boldsymbol{x}_{\mu},\boldsymbol{x}_{\mathrm{pa}(\mu)})$
$q_{-\mu}(\boldsymbol{x}_{-\mu})$ Distribution over states of all gates in circuit not including $\mu$ and $\mathrm{pa}(\mu)$
$\mathcal{B}$ Basis of logic gates
$q_{g}$ Short hand for prior distribution associated with logic gate $g\in\mathcal{B}$
$\tilde{\mathcal{B}}$ Set of $(g,q_{g})$ for every gates $g\in\mathcal{B}$
$q_{g}^{\min}$ Probability of least likely state under distribution $q_{g}$
$\mathrm{I}_{\mu}(\mathrm{X}_{\mu};\mathrm{X}_{\mathrm{ch}(\mu)})$ Mutual information between gate $\mu$ and its children $\mathrm{ch}(\mu)$ with reference to distribution $p^{\mu}(\boldsymbol{x})$
$\mathcal{MC}_{\mu}$ Mismatch cost of running gate $\mu$ in a circuit
$\mathcal{MC}_{l}$ Mismatch cost of running layer $l$ in a circuit
$\mathcal{MC}_{ow}(p_{in})$ Mismatch cost of overwriting inputs with input distribution $p_{in}$
$\mathcal{MC}(C_{n},p_{in})$ Total mismatch cost of running the circuit $C_{n}$ for the input distribution $p_{in}$
Table 1: Table of notation for the main symbols used in the paper.
III.1.1 Notation
We use $p^{\mu}(\boldsymbol{x})$ to denote the distribution over the joint state of the circuit before the execution of gate $\mu$ and after the execution of gate $\mu-1$. Moreover, $p^{0}(\boldsymbol{x})$ denotes the distribution before the over-writing of new input values and $p^{1}(\boldsymbol{x})$ denotes the distribution after the over-writing of new input values but before gate $\mu=1$ is executed.
For any given subset of nodes $\chi\subset V$, we define $X_{\chi}$ as the collection of random variables at the subset of nodes $\chi$, and $\boldsymbol{x}_{\chi}$ as a particular set of values thereof; we denote $p_{\chi}(\boldsymbol{x}_{\chi})$ as the marginal distribution of $X_{\chi}$, obtained from the maximally correlated joint distribution $p^{0}(\boldsymbol{x})$.
$p_{\chi}(\boldsymbol{x}_{\chi})=\sum_{\boldsymbol{x}_{V/\chi}}p^{0}(\boldsymbol{x})$ (7)
For example, we will use $\boldsymbol{x}_{:\mu}$ to denote the joint state of all nodes preceding gate $\mu$, $\boldsymbol{x}_{\mu}$ to denote the state of gate $\mu$ and $\boldsymbol{x}_{\mu:}$ to denote the set of all nodes following gate $\mu$. Accordingly, $p_{:\mu}$ denotes the marginal distribution over the joint state $\boldsymbol{x}_{:\mu}$ of gates preceding the gate $\mu$ obtained from $p^{0}(\boldsymbol{x})$.
$p_{:\mu}(\boldsymbol{x}_{:\mu})=\sum_{\boldsymbol{x}_{\mu},\boldsymbol{x}_{\mu:}}p^{0}(\boldsymbol{x})$ (8)
Analogously, $p_{in}$ and $p_{nin}$ are used to denote the joint distribution over the states of input nodes and non-input gates respectively.
$p_{nin}(\boldsymbol{x}_{nin})=\sum_{x_{in}}p^{0}(\boldsymbol{x})$ (9)
III.1.2 Joint distribution after the re-initialization of the input nodes
As mentioned earlier, when the new inputs are initialized for the next run, the joint state of the non-input gates becomes independent of the new joint state of the input nodes (see Fig. 2B and Fig. 3B). Therefore, after the re-initialization of the input nodes, the joint distribution changes from $p^{0}(\boldsymbol{x})$,
$p^{1}(\boldsymbol{x})=p_{in}(\boldsymbol{x}_{in})p(\boldsymbol{x}_{nin}).$ (10)
Following this re-initialization, the non-input gates run sequentially and the distribution over the joint state evolves accordingly.
III.1.3 Joint distribution after the update of gate $\mu$
As mentioned earlier and depicted in Fig. 2, before gate $\mu$ updates, $\boldsymbol{x}_{\mu}$ is correlated with $\boldsymbol{x}_{\mu:}$ but independent of $\boldsymbol{x}_{:\mu}$. Therefore, the joint distribution before the update of gate $\mu$ is
$p^{\mu}(\boldsymbol{x})=p_{:\mu}(\boldsymbol{x}_{:\mu})p_{\mu,\mu:}(\boldsymbol{x}_{\mu},\boldsymbol{x}_{\mu:}).$ (11)
After the update of the gate $\mu$ based on the new values of its parent gates, $\boldsymbol{x}_{\mu}$ becomes correlated with $\boldsymbol{x}_{:\mu}$ and independent with $\boldsymbol{x}_{\mu:}$ (see Fig. 3). Therefore,
$p^{\mu+1}(\boldsymbol{x})=p_{:\mu,\mu}(\boldsymbol{x}_{:\mu},\boldsymbol{x}_{\mu})p_{\mu:}(\boldsymbol{x}_{\mu:}).$ (12)
$p(\boldsymbol{x}_{:\mu})p(\boldsymbol{x}_{\mu},\boldsymbol{x}_{\mu:})\to p(\boldsymbol{x}_{:\mu},\boldsymbol{x}_{\mu})p(\boldsymbol{x}_{\mu:})$ (13)
III.2 Mismatch cost of re-initializing inputs with new values.
Figure 3: Evolution of joint distribution under re-initialization of inputs and re-running the circuit gate-by-gate. (A) After a complete run, the joint distribution reflects maximal correlation among all gates. This correlation arises from the initial input distribution and the network of dependencies within the circuit. (B) In the next run, the input nodes are re-sampled from the input distribution $p(\boldsymbol{x}_{in})$, making the new inputs independent of the states of non-input gates from the previous run. As a result, the joint distribution after re-initialization is factored as $p(\boldsymbol{x}_{in})p(x_{1},...,x_{5})$, indicating that inputs and non-input gates are now statistically independent. (C) When gate $x_{1}$ updates based on the newly re-sampled input values, it becomes correlated with the input nodes while simultaneously losing correlation with the rest of the gates. Consequently, the joint distribution evolves to $p(\boldsymbol{x}_{in},x_{1})p(x_{2},...,x_{5})$. Similarly, (D), (E), (F), and (G) illustrate the sequential updates of the remaining gates and the corresponding evolution of the distribution. In particular, (G) demonstrates that after a complete run of the circuit, the distribution returns to its initial form as shown in (A).
During the process of overwriting new input values, the input nodes evolve independently of the rest of the non-input nodes which remain unchanged. Therefore, updating input nodes with new values is a sub-system process where the sub-system $\boldsymbol{x}_{in}$ evolves independently of $\boldsymbol{x}_{nin}$ while the later remains unchanged. The prior distribution for this subsystem process is product distribution, expressed as $q_{0}(\boldsymbol{x}_{in})q_{0}(\boldsymbol{x}_{nin})$. Since new values of the input nodes are sampled from $p_{in}(x_{in})$, this prior distribution evolves to $p_{in}(\boldsymbol{x}_{in})q_{0}(\boldsymbol{x}_{nin})$.
$q_{in}(\boldsymbol{x}_{in})q_{nin}(\boldsymbol{x}_{nin})\longrightarrow p_{in}(\boldsymbol{x}_{in})q_{nin}(\boldsymbol{x}_{nin})$ (14)
while the actual distribution evolves from $p^{0}(\boldsymbol{x})$ to $p_{in}(\boldsymbol{x}_{in})p_{nin}(\boldsymbol{x}_{nin})$:
$p^{0}(\boldsymbol{x})\longrightarrow p_{in}(\boldsymbol{x}_{in})p_{nin}(\boldsymbol{x}_{nin})$ (15)
Since new input values are totally independent of the states of the rest of the gates, $\mathrm{I}_{1}(\boldsymbol{x}_{in};\boldsymbol{x}_{nin})=0$, and the drop in mutual information is
$\Delta\mathrm{I}(\mathrm{X}_{in};\mathrm{X}_{nin})$ $=\mathrm{I}_{0}(\mathrm{X}_{in};\mathrm{X}_{nin})-\mathrm{I}_{1}(\mathrm{X}_{in};\mathrm{X}_{nin})$ (16)
$=\mathrm{I}_{0}(\mathrm{X}_{in};\mathrm{X}_{nin})$ (17)
The mismatch cost of overwriting the input values is
$\mathcal{MC}_{ow}=$ $D(p^{0}||q^{0})-D(p^{1}||\tilde{q}^{0})$ (18)
$=$ $\mathrm{I}_{0}(\mathrm{X}_{in};\mathrm{X}_{nin})+D(p_{in}||q_{in})$ (19)
Moreover, if $\mathrm{X}_{1}$ is a deterministic function of $\mathrm{X}_{in}$, then $\mathrm{I}_{0}(\mathrm{X}_{in};\mathrm{X}_{1})=S(p_{in})$. In that case, the overwriting mismatch cost is given by
$\mathcal{MC}_{ow}=\mathcal{C}~{}(p_{in},q_{in})$ (20)
where $\mathcal{C}~{}(p_{in},q_{in})$ of $q_{in}$ relative to $p_{in}$.
III.3 Mismatch cost of gate-by-gate or layer-by-layer implementation of the circuit.
We use $q^{\mu}(\boldsymbol{x})$ for the prior distributions associated with the process of updating the gate $\mu$. Additionally, $q^{0}(\boldsymbol{x})$ is used to denote the prior associated with the process of overwriting. When $\boldsymbol{x}_{\mu}$ is updated based on the values of its parents $\boldsymbol{x}_{\mathrm{pa}(\mu)}$, the rest of the nodes are unchanged. This makes it a subsystem process where $\boldsymbol{x}_{\mu}$ and $\boldsymbol{x}_{\mathrm{pa}(\mu)}$ form a subsystem. The prior distribution therefore is a product distribution of the form
$q^{\mu}(\boldsymbol{x})=q_{\mu,\mathrm{pa}(\mu)}(\boldsymbol{x}_{\mu},\boldsymbol{x}_{\mathrm{pa}(\mu)})q_{-\mu}(\boldsymbol{x}_{-\mu})$ (21)
When gate $\mu$ changes its state based on its parents states, the associated prior distribution $q_{\mu,\mathrm{pa}(\mu)}(\boldsymbol{x}_{\mu},\boldsymbol{x}_{\mathrm{pa}(\mu)})$ evolves to $\tilde{q}_{\mu,\mathrm{pa}(\mu)}(\boldsymbol{x}_{\mu},\boldsymbol{x}_{pa(\mu)})=\pi_{\mu}(\boldsymbol{x}_{\mu}|\boldsymbol{x}_{\mathrm{pa}(\mu)})q_{\mathrm{pa}(\mu)}(\boldsymbol{x}_{\mathrm{pa}(\mu)})$, where $q_{\mathrm{pa}(\mu)}(\boldsymbol{x}_{\mathrm{pa}(\mu)})$ is the marginal distribution of $\mathrm{pa}(\mu)$ obtained from $q_{\mu,\mathrm{pa}(\mu)}(\boldsymbol{x}_{\mu},\boldsymbol{x}_{\mathrm{pa}(\mu)})$. Since the rest of the gates do not change when gate $\mu$ is updated, the distribution $q^{\mu}(\boldsymbol{x})$ evolves to $\tilde{q}^{\mu}(\boldsymbol{x})$ , given by:
$\tilde{q}^{\mu}(\boldsymbol{x})=\tilde{q}_{\mu,\mathrm{pa}(\mu)}(\boldsymbol{x}_{\mu},\boldsymbol{x}_{pa(\mu)})q_{-\mu}(\boldsymbol{x}_{-\mu}).$ (22)
Figure 4: Layer-by-layer implementation of a circuit. (A) State of the circuit after a complete run. (B) Updating of input nodes for the next run, with values re-sampled from $p(\boldsymbol{x}_{in})$. (C) The first layer consists of gates $x_{1}$ and $x_{2}$. As they update together based on the newly sampled input values, they become correlated with the input nodes while losing correlation with the rest of the gates. Consequently, the joint distribution evolves to $p(\boldsymbol{x}_{in},x_{1},x_{2})p(x_{3},x_{4},x_{5})$. (D) Gates $x_{3}$ and $x_{4}$ make up layer 2. As they update based on the new values of gates $x_{1}$ and $x_{2}$, they become correlated with them and the input nodes, while becoming independent of the state of $x_{5}$. The distribution then evolves to $p(\boldsymbol{x}_{in},x_{1},...,x_{4})p(x_{5})$. Finally, in (E), gate $x_{5}$, constituting layer 3, is updated, and the distribution returns to its initial form as shown in (A).
The mismatch cost of updating the gate $\mu$ is given by,
$\mathcal{MC}_{\mu}=D(p^{\mu}||q^{\mu})-D(p^{\mu+1}||\tilde{q}^{\mu})$ (23)
Using equations (11), (12), (21), and (22), the expression for the mismatch cost $\mathcal{MC}_{\mu}$ simplifies to
$\mathcal{MC}_{\mu}=\mathrm{I}_{\mu}(\mathrm{X}_{\mu};\mathrm{X}_{\mathrm{ch}(\mu)})+D\left(p_{\mu}p_{\mathrm{pa}(\mu)}||q_{\mu,\mathrm{pa}(\mu)})\right)-D\left(p_{\mathrm{pa}(\mu)})||q_{\mathrm{pa}(\mu)}\right),$ (24)
where $\mathrm{I}_{\mu}(\mathrm{X}_{\mu};\mathrm{X}_{\mathrm{ch}(\mu)})$ denotes the mutual information between gate $\mu$ and its children $\mathrm{ch}(\mu)$ prior to the update of $\mu$. Specifically, the mutual information is evaluated with respect to the distribution $p^{\mu}(\boldsymbol{x})$, which characterizes the state of the system immediately before the update of $\mu$. The derivation of Eq. (24) is provided in App. A.3. The total mismatch cost of executing the entire circuit through a sequential gate-by-gate implementation is given by the sum of the mismatch costs of individual gates and the mismatch costs incurred when updating input nodes,
$\mathcal{MC}(C_{n},p_{in})=$ $\;\mathcal{MC}_{ow}+\sum_{\mu\in V\backslash V_{in}}\mathcal{MC}_{\mu}$ (25)
$=$ $\;\mathcal{C}~{}(p_{in},q_{in})+\sum_{\mu\in V\backslash V_{in}}\mathrm{I}_{\mu}(\mathrm{X}_{\mu};\mathrm{X}_{\mathrm{ch}(\mu)})+\Delta_{\mu}.$ (26)
where
$\Delta_{\mu}:=D\left(p_{\mu}p_{\mathrm{pa}(\mu)}||q_{\mu,\mathrm{pa}(\mu)})\right)-D\left(p_{\mathrm{pa}(\mu)}||q_{\mathrm{pa}(\mu)}\right).$ (27)
It is straightforward to extend the analysis for a layer-by-layer implementation of circuit (see Fig. 4). Based on a topological ordering of a DAG, it is possible to stratify the set of gates in a natural way into layers. Starting with the set of input nodes $V_{0}:=\{\mu\in V:\mathrm{pa}(\mu)=\emptyset\}$, the set of gates constituting the $l$-th layer is defined as,
$V_{l}:=\{\mu\in V\backslash(V_{0}\cup...\cup V_{l-1}):\mathrm{pa}(\mu)\cap(V_{1}\cup...\cup V_{l-1})\neq\emptyset\},$ (28)
for any $l\in\{1,...,L\}$. Let $X_{l}$ be the random variable denoting the joint state of gates in layer $l$ and let $\boldsymbol{x}_{l}$ denote a value of $X_{l}$. In a layer-by-layer implementation, all the gates in a layer are updated simultaneously. Then, analogous to Eq. 24, the mismatch cost of updating layer $l$ in the circuit is given by,
$\mathcal{MC}_{l}=$ $\mathrm{I}_{l}(\mathrm{X}_{l};\mathrm{X}_{\mathrm{ch}(l)})$ (29)
$+D\left(p_{l}p_{\mathrm{pa}(l)}||q_{l,\mathrm{pa}(l)}\right)-D\left(p_{\mathrm{pa}(l)}||q_{\mathrm{pa}(l)}\right),$
where $p_{l}$, $p_{\mathrm{pa}(l)}$, $q_{l,\mathrm{pa}(l)}$, and $q_{\mathrm{pa}(l)}$ are defined for layer $l$ and have their usual meaning as in Eq. 24. The derivation of Eq. 29 is provided in App. A.3.
Updating multiple gates simultaneously in a layer-by-layer approach, instead of updating gates sequentially one by one, represents a form of time-coarse graining. This is because grouping several updates into a single time step effectively reduces the temporal resolution of the system’s dynamics. As demonstrated in yadav2024mismatch , mismatch cost decreases under time-coarse graining. Applied to circuits, this result implies that the mismatch cost of a layer-by-layer implementation always serves as a lower bound on the mismatch cost of a gate-by-gate implementation. In In Fig. 5, we compare the mismatch costs of the layer-by-layer and gate-by-gate implementations of the same ripple-carry adder. The gate-by-gate mismatch cost is computed using Eq. 24, while the layer-by-layer mismatch cost is computed using Eq. 29. As a consequence of the lowering in mismatch cost under time-coarse graining, the layer-by-layer mismatch cost is lower than the gate-by-gate mismatch cost for the same circuit, across ripple-carry adder circuits with varying input lengths.
Additionally, note that as the computation in a circuit progress—whether gate-by-gate or layer-by-layer—the distribution over the states cycles through a sequence of distributions (see Fig. 3 and 4). It is important to emphasize that, for the calculation of the total mismatch cost, when calculating the total mismatch cost, the phase at which the sequence of distributions is started does not impact the final mismatch cost, as long as the full cycle of execution is completed.
In the next section, we use Eq 24 to calculate the mismatch cost of various circuit families and broadly investigate how does the circuit structure affect their mismatch cost. In particular, we define mismatch cost complexity, analogous to the size and depth complexity of a circuit family. We will then establish results that relate mismatch cost complexity to size complexity.
Figure 5: Comparison of layer-by-layer and gate-by-gate mismatch costs for the ripple-carry adder circuit family. As the circuit size increases with the input size, the mismatch cost grows for both the layer-by-layer (dashed red) and gate-by-gate (solid blue) implementations. Notably, the layer-by-layer mismatch cost remains consistently lower than the gate-by-gate mismatch cost. In this comparison, the input distribution is uniformly random for both implementations, and the priors of all gates in the circuits are also uniform.
IV The mismatch cost complexity of a circuit family
We aim to explore the relationship between the total mismatch cost of a circuit and other complexity measures, such as size and depth. To this end, we extend the definition of a basis for a circuit family to include the prior distribution associated with each logic gate in the basis. For a given basis $\mathcal{B}$, we define the associated prior basis ~=$\tilde{\mathcal{B}}=\{(g,q_{g,\mathrm{pa}(g)}):g\in\mathcal{B}\}$, where $q_{g,\mathrm{pa}(g)}$ denotes the prior distribution associated with gate $g\in\mathcal{B}$. We sometimes use $q_{g}$ as a shorthand for $q_{g,\mathrm{pa}(g)}$. The definition of extended prior basis allows us to define mismatch cost complexity associated with a circuit family.
Figure 6: Binary tree of AND gates. The size of the grows as $n$ for $n$ input nodes whereas its depth grows as $\log(n)$.
Definition 3 (Mismatch Cost Complexity)
Given a basis $\mathcal{B}$ and an associated prior basis $\tilde{\mathcal{B}}$, a circuit family $\{C_{n}\}_{n\in\mathbb{N}}$ is said to have mismatch cost complexity $m(n)$, where $m:\mathbb{N}\to\mathbb{R}^{+}$, if the associated mismatch cost $\mathcal{MC}(C_{n})$ satisfies $\mathcal{MC}(C_{n})\leq m(n)$ for every $n\in\mathbb{N}$.
Analogously, we define the mismatch cost complexity class of functions:
Definition 4 (Mismatch cost complexity class)
Let $m:\mathbb{N}\to\mathbb{R}^{+}$. For a given associated prior basis $\tilde{\mathcal{B}}$, $\mathbf{MMC}(m)$ is the class of functions $f$ for which there exist a circuit family $\mathcal{C}~{}$ of mismatch cost complexity $m$.
The mismatch cost complexity associated with any circuit family is influenced by several key properties of the circuit. First, it depends on the basis $\tilde{\mathcal{B}}$, which includes the prior associated with each allowed logic gate in the circuit. Second, the topology of the circuit, in combination with the input distribution, jointly determines the mutual information $\mathrm{I}_{\mu}(\mathrm{X}_{\mu};\mathrm{X}_{\mathrm{ch}(\mu)})$, which contributes to the mismatch cost for each gate $\mu$ downstream (see Eq. 24). Additionally, the topology and input distribution together determine the joint distribution $p_{\mu}(\boldsymbol{x}_{\mu})$ and $p_{\mu}(\boldsymbol{x}_{\mathrm{pa}}(\mu))$ for each gate $\mu$ downstream.
With these definitions we can derive the following result:
Theorem 1
The mismatch cost of running a gate $\mu$ in a circuit is upper bounded by
$\mathcal{MC}_{\mu}\leq\ln\left(\frac{1}{q_{\mu}^{\min}}\right)$ (30)
where $q_{\mu}^{\min}=\min_{x}q_{\mu}(x)$, regardless of the fan-in or fan-out of the gate, and regardless of the input distribution to the gate.
Moreover, if $q_{\mu}$ is identical for all gates in all circuits in the circuit family, and $p_{in}^{n}$ is the input distribution for the circuit with input size $n$, then the total mismatch cost of any circuit $C_{n}$ in the circuit family is upper bounded by
[…]
The proof is provided in B.1. We refer to a basis in which all prior distributions are identical for all logic gates as a homogeneous associated prior basis. Additionally, note that the term $S(p^{n}_{in})$ in Eq. 31 is upper-bounded by $n\ln 2$.
Thm. 1 establishes a relationship between the size complexity and the mismatch cost complexity of any circuit family constructed from gates in a homogeneous prior basis. If $s(n)$ denotes the size complexity of the circuit family, then according to Eq. (31), the mismatch cost complexity $m(n)$ of the circuit family is given by:
$m(n)\leq s(n)\ln\left(\frac{1}{q^{\min}}\right)+n\ln 2$ (32)
Figure 7: Saturating the Upper Bound on MMC for the AND Binary Tree Circuit Family (Eq. (33)). When the input is always either all ones or all zeros, then $S(p^{n}_{in})=0$ and the theoretical upper bound is $n\ln\left(1\backslash q^{\min}\right)$. The solid blue line represents the total mismatch cost obtained by using Eq. 26 for an input distribution consisting entirely of either all ones or all zeros, matching exactly with the orange dashed line, which corresponds to the theoretical upper bound given by Eq. (33). In contrast, when the input distribution is a uniform distribution, the actual total mismatch cost of AND binary tree remains consistently lower than the theoretical upper bound: the solid green line depicts the total mismatch cost for a uniform input distribution, while the dashed red line shows its respective theoretical upper bound.
This upper bound is also saturable for certain circuit families with certain input distributions. For example, consider the circuit family consisting of AND binary tree. The number of gates in a binary tree with input $n$ is $|C_{n}|=n$. Therefore, according to Eq. (31), an upper bound on the MMC is
[…]
As shown in Fig. 7, when the input is always either all ones or all zeros, the theoretical upper bound is fully attained by the actual total MMC of the circuit. However, for other input distributions, such as a uniform input distribution, the upper bound is not saturated, and the actual total mismatch cost remains consistently lower than the theoretical upper bound.
According to Eq. 32, if the size complexity $s(n)$ of a circuit family grows as a polynomial function of the input size, then the upper bound on the mismatch cost complexity for any homogeneous prior basis also scales polynomially with the input size. In other words, any Boolean function that can be implemented by a polynomial-size circuit can also be realized by a circuit family whose mismatch cost complexity does not exceed polynomial growth. This observation leads to the following corollary:
Corollary 1.1
For any given basis $\mathcal{B}$ and any associated homogeneous prior basis $\tilde{\mathcal{B}}$,
$\mathbf{P_{/poly}}\subseteq\mathbf{MMC}(\mathrm{poly})$ (34)
However, it remains an open question whether the converse holds—that is, whether any function implementable by a circuit family with a mismatch cost that grows no faster than polynomial can also be realized by a circuit family whose size complexity is polynomially bounded, and thus whether $\mathbf{P_{/poly}}=\mathbf{MMC}(\mathrm{poly})$.
Next, we examine the scenario where the prior distribution is not homogeneous across the logic gates in the basis. For a given basis $\mathcal{B}$, we denote the prior distribution associated with each gate $g\in\mathcal{B}$ by $q_{g}$. The following result tells you how the mismatch cost complexity various with input size for an inhomogeneous prior basis:
Theorem 2
For a given basis $\mathcal{B}$, let $\{C_{n}\}_{n\in N}$ be circuit family and let and let $\tilde{\mathcal{B}}$ the associated prior basis. Then, the mismatch cost of circuit of input size $n$ is upper bounded by,
$\mathcal{MC}(C_{n})\leq\sum_{g\in\mathcal{B}}\#_{g}(n)K_{g}+S(p_{in})$ (35)
where #$\#_{g}(n)$ is the number of gates of type $g$ in $C_{n}$ and $K_{g}=\ln\left(1\backslash q_{g}^{\min}\right).$
The proof is provided in App. B.2. Note that Thm. 2 provides a slight generalization of Thm. 1 and applies to heterogeneous prior bases. Several key aspects distinguish it from the homogeneous prior case. For instance, in the homogeneous prior setting, the upper bound on the mismatch cost (MMC) is always proportional to the size complexity $|C_{n}|$ of the circuit family for every $n$. However, with heterogeneous priors, this proportionality may no longer hold.
To illustrate this, consider a circuit family $\{C_{n}\}_{n\in\mathbb{N}}$ whose basis comprises two types of gates — say, AND and XOR — labeled as $1$ and $2$. In this example, the number of AND gates scales as #$\#_{1}(n)=\log(n)$, while the number of XOR gates scales as #$\#_{2}(n)=n^{3}$. According to Thm. 2, the upper bound on the mismatch cost is given by:
$\sum_{g\in\mathcal{B}}\#_{g}(n)K_{g}+S(p^{n}_{in})$ =#$=\#_{1}(n)K_{1}+\#_{2}(n)K_{2}+S(p^{n}_{in})$
$=\log(n)K_{1}+n^{3}K_{2}+S(p^{n}_{in})$ (36)
Now, suppose the priors associated with these two types of gates are such that $K_{1}\gg K_{2}$. Under this condition, for small $n$, the upper bound on the mismatch cost scales as $\log(n)$. However, as $n$ grows large, $n^{3}$ becomes the dominant term. This example demonstrates how, in the case of heterogeneous priors, the MMC can scale differently from the size complexity of the circuit family.
Figure 8: Mismatch cost scaling for a circuit family with homogeneous vs. heterogeneous priors. The circuit family that we consider here consist of binary tree made up AND and OR gates; each layer had one OR gate and rest of the gates are AND. The solid blue curve shows the mismatch cost when the prior of OR gates differs from the uniform prior of AND gates (heterogeneous priors). The mismatch cost is not linear with circuit size (Thm. 2). In contrast, the dashed orange curve represents the mismatch cost for the same circuit family with identical priors for XOR and AND gates (homogeneous priors), and the mismatch cost scales linearly with the circuit size.
In Fig. 8, we demonstrate how, due to the effect of heterogeneous prior distributions, the mismatch cost can deviate from the linear scaling with size.
The lower bound on the MMC
Going back to Eq. (24) for the mismatch cost of a gate in a circuit, under the assumption that the prior distribution $q_{\mu,\mathrm{pa}(\mu)}(\boldsymbol{x}_{\mu},\mathrm{pa}(\mu))$ is a product distribution, $q_{\mu}(\boldsymbol{x}_{\mu})q_{\mathrm{pa}(\mu)}(\boldsymbol{x}_{\mathrm{pa}(\mu)})$, the expression for mismatch cost simplifies to:
$\mathcal{MC}_{\mu}=\mathrm{I}_{0}(\mathrm{X}_{\mu};\mathrm{X}_{\mathrm{ch}(\mu)})+D\left(p_{\mu}\|q_{\mu}\right)$ (37)
It turns out that for certain kinds of circuits, the mutual information between the gate and its children is zero, e.g., in an XOR binary tree. Moreover, if the actual distribution $p_{\mu}$ perfectly matches the prior $q_{\mu}$, the KL-divergence term becomes zero. Therefore, it is theoretically possible for the mismatch cost of a gate to approach zero.
This bound can even be achieved for certain types of circuits with specific prior distributions. For example, in an XOR binary tree fed with a uniform input distribution, the distribution $p_{\mu}$ remains a uniform distribution for every downstream XOR gate in the circuit. If all the XOR gates have a uniform prior, $q_{\mu}$, then the actual distribution $p_{\mu}$ perfectly matches with the prior, leading to zero KL divergence and, hence, zero mismatch cost.
However, even if the mismatch cost of each gate in the circuit is zero, the total mismatch cost includes the overwriting mismatch cost, which is lower bounded by the mutual information between input and non-input gates. Therefore, the total associated mismatch cost is lower bounded by the mutual information between the input and non-input gates,
$\mathcal{MC}(C_{n})\geq\mathrm{I}_{0}(\mathrm{X}_{in},\mathrm{X}_{nin})$ (38)
For a deterministic circuit, the input uniquely determines the state of non-input gates. Therefore, the mutual information simplifies to the entropy of the input distribution.
V Applications to Circuit complexity
The results derived so far have many implications for understanding the thermodynamic costs associated with computation in circuits, as well as for the broader concept of circuit complexity. In this section we briefly sketch a few of them.
V.1 MMC of different circuit families implementing the same function family
The problem of computing a Boolean function with a circuit family under resource constraints naturally leads to the challenge of optimizing both circuit size and depth complexity. Since a Boolean function can be computed by multiple circuit families, size and depth serve as key measures of computational efficiency. The goal, therefore, is to identify the most efficient circuit family—one that minimizes both size and depth complexity. While size complexity corresponds to space efficiency and depth complexity to time efficiency, mismatch cost complexity serves as a measure of thermodynamic efficiency—the minimal dissipated heat in a circuit. Thus, when comparing two circuit families computing the same Boolean function, mismatch cost is a key metric for evaluating thermodynamic efficiency.
Figure 9: The MMC of two different circuit families implementing the addition function, which takes two integers in the standard $n$-bit representation and returns their sum in the standard $(n+1)$-bit representation.
For example, the Boolean function $\text{ADD}^{n}$, which computes the sum of two $n$-bit binary numbers as discussed in Sec. II.1, can be implemented by two circuit families: the ripple-carry adder and the carry look-ahead adder. RCA has linear size and depth complexity, while CLA has size complexity of $n\log n$ and logarithmic depth complexity.
In Fig. 9, we compare the mismatch cost of RCA and CLA across various input sizes. The results show a linear scaling of mismatch cost with size complexity for both circuits. However, as input size increases, RCA consistently exhibits a lower mismatch cost than CLA, indicating that despite being slower, RCA is thermodynamically more efficient.
V.2 MMC of a circuit family for heterogeneous priors
The homogeneous prior basis refers to a scenario where all logical gates in a circuit share the same prior distribution. However, this assumption is rarely valid in practice. The physical implementation of a logic gate often depends on its type; for instance, an AND gate typically has different underlying physical processes than an XOR gate. As a result, just like other physical properties of logic gates, the prior distribution naturally varies with gate type.
This heterogeneity has consequences for the thermodynamic cost of a circuit. Fig. 8 illustrates how mismatch cost changes when considering a circuit family composed of OR and AND gates. The figure demonstrates that the mismatch cost is consistently higher when the prior distribution of OR gates differs from that of AND gates across the circuit family.
Fig. 10 illustrates how the mismatch cost of various individual logic gates changes as the prior distribution shifts. Notably, the mismatch cost increases as the prior distribution deviates further from the uniform distribution. Similarly, Fig. 11 shows how the total mismatch cost (MMC) of a circuit composed of AND and OR gates varies as the prior distributions of the two gate types become increasingly heterogeneous.
Figure 10: Mismatch cost of individual gates as a function of their prior distributions. The input distribution to the gate is i.i.d uniform.
V.3 Smallest MMC and smallest size circuit implementing a given function
Finding the smallest-size circuit for a given Boolean function is a central problem in circuit complexity and often involves optimizing for minimal redundancy. However, the circuit with the smallest size does not necessarily minimize total MMC. For instance, consider computing the NAND function: while it can be implemented with a single NAND gate, an alternative circuit using an AND gate followed by a NOT gate might result in a lower MMC if the NAND gate’s prior is highly non-uniform, while the priors for AND and NOT are closer to uniform. This example illustrates that optimizing for thermodynamic cost may require different design choices than optimizing for size alone, and raises the broader question of trade-offs between logical and energetic efficiency in circuit design.
Figure 11: Total mismatch cost of a circuit with 4 layers, each containing two AND gates and one OR gate. The prior distribution for AND gates is fixed and uniform, while the prior of OR gates is varied. The plot shows how the total mismatch cost changes with the OR gate prior.
VI Discussion and Future work
Our work establishes a framework for analyzing the unavoidable thermodynamic costs of computation in Boolean circuits by deriving the expression for mismatch cost. This formulation allows for a direct comparison between circuit families based on their mismatch cost complexity, providing a new perspective on circuit optimization—one that extends beyond minimizing size and depth to include energetic efficiency.
A key result of our study is Th. 1, which establishes an upper bound on the mismatch cost complexity that scales linearly with the size complexity of a circuit family under the assumption of a homogeneous prior distribution across all gates. Understanding precisely what causes the slope differences between circuit families, is still a question. In contrast to Th. 1, Th. 2 relaxes this assumption, allowing for cases where mismatch cost complexity deviates from size complexity. This result opens avenues for further exploration into how circuit structure, gate heterogeneity, and prior distributions influence thermodynamic costs. For example, with known prior distributions for each type of gate, a key question arises: how can we optimize the circuit topology associated with a Boolean function to minimize the thermodynamic cost?
Another open question concerns the role of fan-in and fan-out—the number of inputs to, and outputs from, a gate, respectively—in determining the total mismatch cost. Fan-out greater than one allows the same gate output to be reused in multiple locations, often reducing the overall circuit size. However, it remains unclear how such reuse affects the minimal thermodynamic cost. Does increasing fan-out reduce the mismatch cost, just as it reduces size? Or could it increase or leave the cost unchanged? Investigating this trade-off between logical reuse and energetic dissipation is an important direction for future work.
A fundamental problem is identifying the circuit family for a given Boolean function that minimizes mismatch cost. Understanding how this circuit family differs from those optimized for size or depth complexity would clarify the fundamental trade-offs between space, time, and thermodynamic cost in computation. Another intriguing direction is the thermodynamic analysis of circuit classes such as NC, which consists of functions that can be efficiently parallelized. A key open question is whether functions in NC—those that can be computed efficiently in parallel—are also implementable by circuits with low thermodynamic cost. Investigating this could provide new insights into the interplay between parallelizability and energetic efficiency in computation.
VII Acknowledgement
This work was supported by the U.S. National Science Foundation (NSF) Grant 2221345. We thank Santa Fe Institute for helping to support this research. We thank Harrison Hartle for valuable discussions and insightful feedback during the development of this work.
Appendix A Derivation of Eq. 24
A.1 Subsystem process
Consider a system composed of two subsystems $A$ and $B$ with state space $\mathcal{X}_{\mu}$ and $\mathcal{X}_{B}$ respectively. A process that evolves initial distribution $p^{0}_{AB}(x_{A},x_{B})$ over the joint state space $\mathcal{X}_{\mu}\times\mathcal{X}_{B}$ is called a subsystem process if the two subsystem evolve independent of each other during the process
$G_{AB}(x^{\prime}_{A},x^{\prime}_{B}|x_{A},x_{B})=G_{A}(x^{\prime}_{A}|x_{A})G_{B}(x^{\prime}_{B}|x_{B})$ (39)
and entropy flow of the joint system is the sum of the entropy flow of the subsystems,
$Q(p^{0}_{AB})=Q_{A}(p^{0}_{A})+Q_{B}(p^{0}_{B})$ (40)
where $Q_{A}(p^{0}_{A})$ and $Q_{B}(p^{0}_{B})$ can be referred as subsystem EF.
If $q^{0}(x_{A})$ and $q^{0}(x_{B})$ are the prior distributions for the evolution of subsystems $A$ and $B$ respectively, the prior distribution for the joint evolution of A and B, such that 39 and 40 hold, is the product distribution,
$q^{0}_{AB}(x_{A},x_{B})=q^{0}_{A}(x_{A})q^{0}_{B}(x_{B})$ (41)
A.2 Mismatch Cost lower bound on EP of a subsystem process
Consider a system $S$ consisting of multiple subsystems that evolve one after another, and the dynamics of one or more of these subsystems does not directly depend on the state of some other subsystems. In particular, let us focus on a subsystem $\mu$ that evolves based on the value of another subsystem denoted as $\mathrm{pa}(\mu)$. The set of the rest of the subsystems is denoted as $\tilde{\mu}$. This results in a particular partition of the set of subsystems, $S=\mu\cup\mathrm{pa}(\mu)\cup\tilde{\mu}$. Let us denote $p^{0}$ the joint probability distribution of the system before subsystem $\mu$ updates. Accordingly, let $p^{1}$ denote the joint probability distribution after subsystem $\mu$ updates, while the rest of the system remains the same. The distribution undergoes a change,
$p^{0}(\boldsymbol{x}_{\mu},\boldsymbol{x}_{\mathrm{pa}(\mu)},\boldsymbol{x}_{\tilde{\mu}})\longrightarrow p^{1}(\boldsymbol{x}_{\mu},\boldsymbol{x}_{\mathrm{pa}(\mu)},\boldsymbol{x}_{\tilde{\mu}}),$ (42)
such that $p^{1}_{\mu,\mathrm{pa}(\mu)}(\boldsymbol{x}_{\mu},\boldsymbol{x}_{\mathrm{pa}(\mu)})=\sum_{\boldsymbol{x}_{\tilde{\mu}}}p^{1}(\boldsymbol{x})=\pi_{\mu}(\boldsymbol{x}_{\mu}|\boldsymbol{x}_{\mathrm{pa}(\mu)})p^{0}_{\mu}(\boldsymbol{x}_{\mu})$, and since the rest of the subsystems do not change, we have $p^{1}_{\tilde{\mu}}(\boldsymbol{x}_{\tilde{\mu}})=\sum_{\boldsymbol{x}_{\mu},\boldsymbol{x}_{\mathrm{pa}(\mu)}}p^{1}(\boldsymbol{x})=p^{0}_{\tilde{\mu}}(\boldsymbol{x}_{\tilde{\mu}})$.
As mentioned above in A.1, the prior distribution for this kind of subsystem process is a product distribution of the form $q^{0}(\boldsymbol{x})=q^{0}(\boldsymbol{x}_{\mu},\boldsymbol{x}_{\mathrm{pa}(\mu)}))q^{0}(\boldsymbol{x}_{\tilde{\mu}})$. Under the evolution of the subsystem $\mu$ conditioned on the state of subsystem $\mathrm{pa}(\mu)$, the prior distribution evolves to $q^{1}(\boldsymbol{x})=q^{1}(\boldsymbol{x}_{\mu},\boldsymbol{x}_{\mathrm{pa}(\mu)})q^{0}(\boldsymbol{x}_{\tilde{\mu}})$.
$q^{0}(\boldsymbol{x}_{\mu},\boldsymbol{x}_{\mathrm{pa}(\mu)}))q^{0}(\boldsymbol{x}_{\tilde{\mu}})\longrightarrow q^{1}(\boldsymbol{x}_{\mu},\boldsymbol{x}_{\mathrm{pa}(\mu)})q^{0}(\boldsymbol{x}_{\tilde{\mu}})$ (43)
where $q^{1}_{\mu,\mathrm{pa}(\mu)}(\boldsymbol{x}_{\mu},\boldsymbol{x}_{\mathrm{pa}(\mu)})=\pi_{\mu}(\boldsymbol{x}_{\mu}|\boldsymbol{x}_{\mathrm{pa}(\mu)})q^{0}_{\mathrm{pa}(\mu)}(\boldsymbol{x}_{\mathrm{pa}(\mu)})$.
The mismatch cost of the subsystem process is given by,
$\mathcal{MC}=D\left(p^{0}\|q^{0}\right)-D\left(p^{1}\|q^{1}\right)$ (44)
where the first KL divergence can be expanded to,
$D\left(p^{0}\|q^{0}\right)=\mathrm{I}_{0}\left(\mathrm{X}_{\mu},\mathrm{X}_{\mathrm{pa}(\mu)};\mathrm{X}_{\tilde{\mu}}\right)+D\left(p^{0}_{\mu,\mathrm{pa}(\mu)}\|q^{0}_{\mu,\mathrm{pa}(\mu)}\right)+D\left(p^{0}_{\tilde{\mu}}\|q^{0}_{\tilde{\mu}}\right),$ (45)
and the second KL divergence,
$D\left(p^{1}\|q^{1}\right)=\mathrm{I}_{1}\left(\mathrm{X}_{\mu},\mathrm{X}_{\mathrm{pa}(\mu)};\mathrm{X}_{\tilde{\mu}}\right)+D\left(p^{1}_{\mu,\mathrm{pa}(\mu)}\|q^{1}_{\mu,\mathrm{pa}(\mu)}\right)+D\left(p^{0}_{\tilde{\mu}}\|q^{0}_{\tilde{\mu}}\right).$ (46)
Since $p^{1}_{\mu,\mathrm{pa}(\mu)}(\boldsymbol{x}_{\mu},\boldsymbol{x}_{\mathrm{pa}(\mu)})=\pi_{\mu}(\boldsymbol{x}_{\mu}|\boldsymbol{x}_{\mathrm{pa}(\mu)})p^{0}_{\mathrm{pa}(\mu)}(\boldsymbol{x}_{\mathrm{pa}(\mu)})$ and $q^{1}_{\mu,\mathrm{pa}(\mu)}(\boldsymbol{x}_{\mu},\boldsymbol{x}_{\mathrm{pa}(\mu)})=\pi_{\mu}(\boldsymbol{x}_{\mu}|\boldsymbol{x}_{\mathrm{pa}(\mu)})q^{0}_{\mathrm{pa}(\mu)}(\boldsymbol{x}_{\mathrm{pa}(\mu)})$, therefore, using chain rule for KL divergence, $D\left(p^{1}_{\mu,\mathrm{pa}(\mu)}\|q^{1}_{\mu,\mathrm{pa}(\mu)}\right)=D\left(p^{0}_{\mathrm{pa}(\mu)}\|q^{0}_{\mathrm{pa}(\mu)}\right)$. Therefore,
$D(p^{1}\|q^{1})=\mathrm{I}_{1}\left(\mathrm{X}_{\mu},\mathrm{X}_{\mathrm{pa}(\mu)};\mathrm{X}_{\tilde{\mu}}\right)+D\left(p^{0}_{\mathrm{pa}(\mu)}\|q^{0}_{\mathrm{pa}(\mu)}\right)+D\left(p^{0}_{\tilde{\mu}}\|q^{0}_{\tilde{\mu}}\right)$ (47)
Combining equation (45) and (47), the mismatch cost can be written as,
$\mathcal{MC}=\Delta\mathrm{I}+D\left(p^{0}_{\mu,\mathrm{pa}(\mu)}\|q^{0}_{\mu,\mathrm{pa}(\mu)}\right)-D\left(p^{0}_{\mathrm{pa}(\mu)}\|q^{0}_{\mathrm{pa}(\mu)}\right)$ (48)
where $\Delta\mathrm{I}=\mathrm{I}_{1}\left(\mathrm{X}_{\mu},\mathrm{X}_{\mathrm{pa}(\mu)};\mathrm{X}_{\tilde{\mu}}\right)-\mathrm{I}_{0}\left(\mathrm{X}_{\mu},\mathrm{X}_{\mathrm{pa}(\mu)};\mathrm{X}_{\tilde{\mu}}\right)$ is the drop in mutual information between the subsystem consisting of $\boldsymbol{x}_{\mu}$ and $\boldsymbol{x}_{\mathrm{pa}(\mu)}$ and the rest of the system.
A.3 Application of subsystem mismatch cost to circuits
When a gate or a layer of gates in a circuit is updated based on the state of the parent gates (or parent layers) while the rest of the circuit remains unchanged, it constitutes a subsystem process. Once again, let’s denote $\mu$ the gate or the layer of gates As described in Sec III.1.3, under the update of gate or layer $\mu$, the actual distribution of the circuit changes from $p^{\mu}(\boldsymbol{x})=p_{:\mu}(\boldsymbol{x}_{:\mu})p_{\mu,\mu:}(\boldsymbol{x}_{\mu},\boldsymbol{x}_{\mu:})$ to $p^{\mu+1}(\boldsymbol{x})=p_{:\mu,\mu}(\boldsymbol{x}_{:\mu},\boldsymbol{x}_{\mu})p_{\mu:}(\boldsymbol{x}_{\mu:})$. The mutual information between the subsystem $\{\mu,\mathrm{pa}(\mu)\}$ and the rest of the circuit can be decomposed into the sum of two components: the mutual information between $\mu$ and its children $\mathrm{ch}(\mu)\subset\tilde{\mu}$, and the mutual information between $\mathrm{pa}(\mu)$ and its directly connected neighborhood of gates, which includes its parent $\mathrm{pa}(\mathrm{pa}(\mu))$ and its children $\mathrm{ch}(\mathrm{pa}(\mu))$,
$\mathrm{I}_{\mu}(\mathrm{X}_{\mu},\mathrm{X}_{\mathrm{pa}(\mu)};\mathrm{X}_{\tilde{\mu}})=\mathrm{I}_{\mu}(\mathrm{X}_{\mu};\mathrm{X}_{\mathrm{ch}(\mu)})+\mathrm{I}_{\mu}(\mathrm{X}_{\mathrm{pa}(\mu)};\mathrm{X}_{\mathrm{pa}(\mathrm{pa}(\mu))},\mathrm{X}_{\mathrm{ch}(\mathrm{pa}(\mu))})$ (49)
In the distribution $p^{\mu+1}(\boldsymbol{x})$, the variables $\mathrm{X}_{\mu}$ and $\mathrm{X}_{\mathrm{ch}(\mu)}$ are independent, implying that $\mathrm{I}_{\mu+1}(\mathrm{X}_{\mu};\mathrm{X}_{\mathrm{ch}(\mu)})=0$. On the other hand, the mutual information between $\mathrm{pa}(\mu)$ and its directly connected neighborhood of gates remains the same in the distribution $p^{\mu+1}$ as in $p^{\mu}$. Therefore,
$\Delta\mathrm{I}$ $=\mathrm{I}_{\mu}(\mathrm{X}_{\mu},\mathrm{X}_{\mathrm{pa}(\mu)};\mathrm{X}_{\tilde{\mu}})-\mathrm{I}_{\mu+1}(\mathrm{X}_{\mu},\mathrm{X}_{\mathrm{pa}(\mu)};\mathrm{X}_{\tilde{\mu}})$ (50)
$=\mathrm{I}_{\mu}(\mathrm{X}_{\mu};\mathrm{X}_{\mathrm{ch}(\mu)}).$ (51)
Using Eq. 48, the mismatch cost of updating gate $\mu$ is,
$\mathcal{MC}_{\mu}=\mathrm{I}_{\mu}(\mathrm{X}_{\mu};\mathrm{X}_{\mathrm{ch}(\mu)})+D\left(p_{\mu}p_{\mathrm{pa}(\mu)}||q_{\mu,\mathrm{pa}(\mu)})\right)-D\left(p_{\mathrm{pa}(\mu)})||q_{\mathrm{pa}(\mu)}\right)$ (52)
The joint distribution of the circuit in a layer-by-layer implementation evolves in a manner similar to that of the gate-by-gate implementation (Fig. 4). The distribution of the circuit before the update of layer $l$ is,
$p^{l}(\boldsymbol{x})=p_{:l}(\boldsymbol{x}_{:l})\,p_{l,l:}(\boldsymbol{x}_{l},\boldsymbol{x}_{l:}).$ (53)
After the update of layer $l$, the joint distribution evolves to,
$p^{l+1}(\boldsymbol{x})=p_{:l,l}(\boldsymbol{x}_{:l},\boldsymbol{x}_{l})p_{l:}(\boldsymbol{x}_{l:}).$ (54)
The prior associated with the update of layer $l$ is defined similar to Eq. (21),
$q^{l}(\boldsymbol{x})=q_{l,\mathrm{pa}(l)}(\boldsymbol{x}_{l},\boldsymbol{x}_{\mathrm{pa}(l)})q_{-1}(\boldsymbol{x}_{-(l\cup\mathrm{pa}(l))})$ (55)
which after the update of layer $l$ evolves to,
$\tilde{q}^{l}(\boldsymbol{x})=\tilde{q}_{l,\mathrm{pa}(l)}(\boldsymbol{x}_{l},\boldsymbol{x}_{\mathrm{pa}(l)})q_{/(l\cup\mathrm{pa}(l))}(\boldsymbol{x}_{/(l\cup\mathrm{pa}(l))})$ (56)
The associated mismatch cost is,
$\mathcal{MC}_{l}$ $=D(p^{l}||q^{l})-D(p^{l+1}||\tilde{q}^{l})$ (57)
$=\mathrm{I}_{l}(\mathrm{X}_{l};\mathrm{X}_{\mathrm{ch}(l)})+D\left(p_{l}p_{\mathrm{pa}(l)}||q_{l,\mathrm{pa}(l)}\right)-D\left(p_{\mathrm{pa}(l)}||q_{\mathrm{pa}(l)}\right)$
Appendix B Proofs
B.1 Proof of Thm. 1
The mismatch cost associated with the execution of a gate $\mu$ in the middle of a run is given by Eq. (24),
$\mathcal{MC}_{\mu}=\mathrm{I}_{0}(\boldsymbol{x}_{\mu};\boldsymbol{x}_{\mathrm{ch}(\mu)})+\Delta_{\mu}$ (58)
where
$\Delta_{\mu}:=D(p_{0}(\boldsymbol{x}_{\mu},\boldsymbol{x}_{\mathrm{pa}(\mu)})||q_{0}(\boldsymbol{x}_{\mu},\boldsymbol{x}_{\mathrm{pa}(\mu)}))-D(p_{0}(\boldsymbol{x}_{\mathrm{pa}(\mu)})||q_{0}(\boldsymbol{x}_{\mathrm{pa}(\mu)})).$
Note that $p_{0}(\boldsymbol{x}_{\mu},\boldsymbol{x}_{\mathrm{pa}(\mu)})$ is the distribution over the joint state of $\mu$ and $\mathrm{pa}(\mu)$ before the execution of gate $\mu$ when $\mu$ and $\mathrm{pa}(\mu)$ are independent of each other. Therefore, $p_{0}(\boldsymbol{x}_{\mu},\boldsymbol{x}_{\mathrm{pa}(\mu)})$ can be written as a product distribution,
$p_{0}(\boldsymbol{x}_{\mu},\boldsymbol{x}_{\mathrm{pa}(\mu)})=p_{0}(\boldsymbol{x}_{\mu})p_{0}(\boldsymbol{x}_{\mathrm{pa}(\mu)})$ (59)
For simplification we assume that $q_{0}$ is a product distribution, i.e., $q_{0}(\boldsymbol{x}_{\mu},\boldsymbol{x}_{\mathrm{pa}(\mu)})=q_{0}(\boldsymbol{x}_{\mu})q_{0}(\boldsymbol{x}_{\mathrm{pa}(\mu)})$. In that case, Eq. 58 simplfies to
$\mathcal{MC}_{\mu}=\mathrm{I}_{\mu}(\mathrm{X}_{\mu};\mathrm{X}_{\mathrm{ch}(\mu)})+D\left(p_{\mu}||q_{\mu}\right).$ (60)
The mutual information between a gate $\mu$ and its children gates $\mathrm{ch}(\mu)$ is always upper bounded by the entropy of the gate $\mu$,
$\mathcal{MC}_{\mu}$ $\leq S(p_{\mu})+D(p_{\mu})\|q_{\mu}))$ (61)
$=\mathcal{C}~{}(p_{\mu},q_{\mu})$ (62)
where $\mathcal{C}~{}(p_{\mu},q_{\mu})$ is the cross entropy of the distribution $p_{\mu}$ relative to a distribution $q_{\mu}$. Moreover, the above inequality is an equality when $\boldsymbol{x}_{\mu}$ is completely determined by $\boldsymbol{x}_{\mathrm{ch}(\mu)}$ ($S(\mathrm{X}_{\mu}|\mathrm{X}_{\mathrm{ch}(\mu)})=0$).
Furthermore, regardless of the actual distribution $p_{\mu}(\boldsymbol{x}_{\mu})$, the cross-entropy is upper bounded by $\ln(1/q_{\mu}^{\min})$. Therefore,
$\mathcal{MC}_{\mu}\leq\ln\left(\frac{1}{q_{\mu}^{\min}}\right)$ (63)
If all the gates in the circuit have the same prior, the sum total of the mismatch cost of running all the gates is upper bounded by,
$\sum_{\mu\in V_{nin}}\mathcal{MC}_{max}=|C_{n}|\ln\left(\frac{1}{q_{0}^{\min}}\right)$ (64)
where $|C_{n}|$ is the number of gates in the circuit. Note that the Eq. 64 does not include the cost of overwriting. So the true total cost is going to be the sum of (64) and the overwriting cost. From Eq. (20), the mismatch cost associated with the overwriting is upper bounded by $S(p_{in})$. Therefore, total associated mismatch cost of a circuit is,
$\mathcal{MC}(C_{n})\leq|C_{n}|\ln\left(\frac{1}{q^{\min}}\right)+S(p_{in})$ (65)
B.2 Proof of Thm. 2
From Eq. 63, the mismatch cost for each gate is upper bound by $\ln(1\backslash q_{\mu}^{\min})$. Let us denote thee prior distribution of logic gate $g\in\mathcal{B}$ as $q_{g}$, and $K_{g}=\ln(1\backslash q_{\mu}^{\min})$. If #$\#_{g}(n)$ is the number of gates of type $g\in\mathcal{B}$ in a circuit $C_{n}$, then the sum total of mismatch cost of gates is upper bounded.
$\sum_{\mu=0}^{|C_{n}|}\mathcal{MC}_{\mu}\leq\sum_{g\in\mathcal{B}}\#_{g}(n)K_{g}$ (66)
Since the overwriting mismatch cost $\mathcal{MC}(p_{i}n)\leq S(p_{in})$, therefore, the total mismatch cost is upper bounded by,
$\mathcal{MC}(C_{n},p_{in})\leq\sum_{g\in\mathcal{B}}\#_{g}(n)K_{g}+S(p_{in})$ (67)
We can also express the summation term as
$\sum_{g\in\mathcal{B}}\#_{g}(n)K_{g}=|C_{n}|\sum_{g\in\mathcal{B}}\gamma_{g}(n)K_{g}$ (68)
where $\gamma_{g}(n)=\#_{g}(n)/|C_{n}|$. Note that if $\gamma_{g}(n)$ is a constant, i.e., $\gamma_{g}(n)=\gamma_{g}$ for all $n\in\mathbb{N}$ (or for $n\to\infty$), then we get the upper bound similar to Th. 1:
$\mathcal{MC}(C_{n})\leq|C_{n}|K^{\prime}+S(p_{in})$ (69)
where $K^{\prime}=\sum_{g\in\mathcal{B}}\gamma_{g}K_{g}$.